# DiffSampling: Enhancing Diversity and Accuracy in Neural Text Generation

### v4

@@ D.4 Ablation Study on the Dynamic Upper Bound
DiffS.-cut $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.60_{\pm.01}$ $0.28_{\pm.00}$ $0.64_{\pm.01}$ $0.26_{\pm.01}$ $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.44_{\pm.01}$ $0.21_{\pm.00}$ $0.46_{\pm.01}$ $0.19_{\pm.01}$
DiffS.-lb $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.54_{\pm.01}$ $0.26_{\pm.00}$ $0.57_{\pm.01}$ $0.24_{\pm.01}$ $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.55_{\pm.01}$ $0.28_{\pm.00}$ $0.60_{\pm.01}$ $0.26_{\pm.01}$
DiffS.-minp $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.53_{\pm.01}$ $0.26_{\pm.00}$ $0.56_{\pm.01}$ $0.24_{\pm.01}$ $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.55_{\pm.01}$ $0.27_{\pm.00}$ $0.59_{\pm.01}$ $0.25_{\pm.01}$
@@
DiffS.-cut $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.49_{\pm.01}$ $0.24_{\pm.00}$ $0.52_{\pm.01}$ $0.22_{\pm.01}$ $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.44_{\pm.01}$ $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.21_{\pm.01}$
DiffS.-lb $0.06_{\pm.00}$ $0.15_{\pm.01}$ $0.28_{\pm.01}$ $0.65_{\pm.01}$ $0.64_{\pm.01}$ $0.90_{\pm.01}$ $0.77_{\pm.01}$ $0.23_{\pm.00}$ $0.51_{\pm.01}$ $0.66_{\pm.01}$ $0.64_{\pm.01}$ $0.37_{\pm.00}$ $0.72_{\pm.01}$ $0.33_{\pm.01}$
DiffS.-minp $0.23_{\pm.00}$ $0.52_{\pm.01}$ $0.66_{\pm.01}$ $0.61_{\pm.01}$ $0.36_{\pm.00}$ $0.72_{\pm.01}$ $0.32_{\pm.01}$ $0.24_{\pm.00}$ $0.53_{\pm.01}$ $0.68_{\pm.00}$ $0.60_{\pm.01}$ $0.33_{\pm.00}$ $0.68_{\pm.01}$ $0.30_{\pm.01}$
@@
DiffS.-cut $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.51_{\pm.01}$ $0.25_{\pm.00}$ $0.56_{\pm.01}$ $0.23_{\pm.01}$ $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.44_{\pm.01}$ $0.22_{\pm.00}$ $0.49_{\pm.01}$ $0.21_{\pm.01}$
DiffS.-lb $01_{\pm.00} 0.0$$03_{\pm.00}$ $0.13_{\pm.00}$ $0.64_{\pm.01}$ $0.63_{\pm.00}$ $0.94_{\pm.00}$ $0.94_{\pm.00}$ $0.22_{\pm.00}$ $0.50_{\pm.01}$ $0.64_{\pm.01}$ $0.61_{\pm.01}$ $0.41_{\pm.00}$ $0.74_{\pm.01}$ $0.35_{\pm.01}$
DiffS.-minp $0.21_{\pm.00}$ $0.49_{\pm.01}$ $0.63_{\pm.01}$ $0.64_{\pm.01}$ $0.44_{\pm.00}$ $0.78_{\pm.01}$ $0.38_{\pm.01}$ $0.24_{\pm.00}$ $0.53_{\pm.01}$ $0.68_{\pm.00}$ $0.63_{\pm.01}$ $0.33_{\pm.00}$ $0.70_{\pm.01}$ $0.30_{\pm.01}$
@@
DiffS.-cut $0.20_{\pm.00}$ $0.45_{\pm.01}$ $0.59_{\pm.01}$ $0.66_{\pm.01}$ $0.42_{\pm.01}$ $0.71_{\pm.01}$ $0.38_{\pm.01}$ $0.25_{\pm.00}$ $0.54_{\pm.01}$ $0.69_{\pm.00}$ $0.45_{\pm.01}$ $0.22_{\pm.00}$ $0.50_{\pm.01}$ $0.21_{\pm.01}$
DiffS.-lb $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.75_{\pm.01}$ $0.61_{\pm.00}$ $0.96_{\pm.00}$ $0.95_{\pm.00}$ $0.20_{\pm.00}$ $0.46_{\pm.01}$ $0.61_{\pm.01}$ $0.65_{\pm.01}$ $0.44_{\pm.01}$ $0.79_{\pm.01}$ $0.40_{\pm.01}$
DiffS.-minp $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.75_{\pm.01}$ $0.61_{\pm.00}$ $0.96_{\pm.00}$ $0.95_{\pm.00}$ $0.23_{\pm.00}$ $0.52_{\pm.01}$ $0.67_{\pm.00}$ $0.62_{\pm.01}$ $0.35_{\pm.00}$ $0.72_{\pm.01}$ $0.32_{\pm.01}$
@@
DiffS.-cut $0.20_{\pm.00}$ $0.49_{\pm.01}$ $0.73_{\pm.01}$ $0.53_{\pm.01}$ $0.31_{\pm.00}$ $0.47_{\pm.01}$ $0.29_{\pm.01}$ $0.20_{\pm.00}$ $0.51_{\pm.00}$ $0.77_{\pm.01}$ $0.39_{\pm.01}$ $0.20_{\pm.00}$ $0.31_{\pm.01}$ $0.18_{\pm.01}$
DiffS.-lb $0.19_{\pm.00}$ $0.47_{\pm.01}$ $0.71_{\pm.01}$ $0.52_{\pm.01}$ $0.39_{\pm.01}$ $0.49_{\pm.01}$ $0.31_{\pm.01}$ $0.19_{\pm.00}$ $0.46_{\pm.01}$ $0.68_{\pm.01}$ $0.56_{\pm.01}$ $0.43_{\pm.01}$ $0.53_{\pm.01}$ $0.34_{\pm.01}$
DiffS.-minp $0.19_{\pm.00}$ $0.48_{\pm.01}$ $0.73_{\pm.01}$ $0.50_{\pm.01}$ $0.35_{\pm.01}$ $0.46_{\pm.01}$ $0.28_{\pm.01}$ $0.19_{\pm.00}$ $0.47_{\pm.01}$ $0.70_{\pm.01}$ $0.53_{\pm.01}$ $0.39_{\pm.01}$ $0.50_{\pm.01}$ $0.32_{\pm.01}$
@@
DiffS.-cut $0.20_{\pm.00}$ $0.50_{\pm.00}$ $0.76_{\pm.01}$ $0.43_{\pm.01}$ $0.24_{\pm.00}$ $0.36_{\pm.01}$ $0.22_{\pm.01}$ $0.20_{\pm.00}$ $0.51_{\pm.00}$ $0.77_{\pm.01}$ $0.40_{\pm.01}$ $0.21_{\pm.00}$ $0.34_{\pm.01}$ $0.19_{\pm.01}$
DiffS.-lb $02_{\pm.00} 0.0$$06_{\pm.00}$ $0.19_{\pm.00}$ $0.81_{\pm.01}$ $0.70_{\pm.00}$ $0.97_{\pm.00}$ $0.90_{\pm.00}$ $0.11_{\pm.00}$ $0.23_{\pm.01}$ $0.34_{\pm.01}$ $0.77_{\pm.01}$ $0.79_{\pm.00}$ $0.89_{\pm.00}$ $0.70_{\pm.01}$
DiffS.-minp $0.16_{\pm.00}$ $0.37_{\pm.01}$ $0.52_{\pm.01}$ $0.78_{\pm.01}$ $0.62_{\pm.00}$ $0.79_{\pm.01}$ $0.53_{\pm.01}$ $0.18_{\pm.00}$ $0.43_{\pm.01}$ $0.61_{\pm.01}$ $0.69_{\pm.01}$ $0.52_{\pm.00}$ $0.68_{\pm.01}$ $0.44_{\pm.01}$
@@
DiffS.-cut $0.20_{\pm.00}$ $0.50_{\pm.00}$ $0.76_{\pm.01}$ $0.46_{\pm.01}$ $0.26_{\pm.00}$ $0.40_{\pm.01}$ $0.24_{\pm.01}$ $0.20_{\pm.00}$ $0.50_{\pm.00}$ $0.77_{\pm.01}$ $0.41_{\pm.01}$ $0.21_{\pm.00}$ $0.34_{\pm.01}$ $0.20_{\pm.01}$
DiffS.-lb $01_{\pm.00} 0.0$$03_{\pm.00}$ $0.15_{\pm.00}$ $0.81_{\pm.01}$ $0.60_{\pm.00}$ $0.98_{\pm.00}$ $94_{\pm.00} 0.0$$08_{\pm.00}$ $0.19_{\pm.01}$ $0.29_{\pm.01}$ $0.78_{\pm.01}$ $0.82_{\pm.00}$ $0.92_{\pm.00}$ $0.76_{\pm.01}$
DiffS.-minp $0.12_{\pm.00}$ $0.29_{\pm.01}$ $0.41_{\pm.01}$ $0.81_{\pm.01}$ $0.73_{\pm.00}$ $0.89_{\pm.00}$ $0.64_{\pm.01}$ $0.18_{\pm.00}$ $0.42_{\pm.01}$ $0.60_{\pm.01}$ $0.70_{\pm.01}$ $0.54_{\pm.00}$ $0.71_{\pm.01}$ $0.46_{\pm.01}$
@@
DiffS.-cut $0.07_{\pm.00}$ $0.17_{\pm.01}$ $0.32_{\pm.01}$ $0.51_{\pm.01}$ $0.57_{\pm.01}$ $0.72_{\pm.01}$ $0.72_{\pm.01}$ $0.20_{\pm.00}$ $0.50_{\pm.00}$ $0.77_{\pm.01}$ $0.41_{\pm.01}$ $0.21_{\pm.00}$ $0.35_{\pm.01}$ $0.20_{\pm.01}$
DiffS.-lb $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.76_{\pm.01}$ $0.60_{\pm.00}$ $0.97_{\pm.00}$ $95_{\pm.00} 0.0$$04_{\pm.00}$ $0.12_{\pm.00}$ $0.24_{\pm.01}$ $0.79_{\pm.01}$ $0.80_{\pm.00}$ $0.95_{\pm.00}$ $0.83_{\pm.00}$
DiffS.-minp $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.76_{\pm.01}$ $0.60_{\pm.00}$ $0.97_{\pm.00}$ $0.95_{\pm.00}$ $0.17_{\pm.00}$ $0.40_{\pm.01}$ $0.56_{\pm.01}$ $0.75_{\pm.01}$ $0.58_{\pm.00}$ $0.77_{\pm.01}$ $0.50_{\pm.01}$

### v5

@@ F.2 Extreme Summarization
Similar considerations can be traced for XSum, as reported in Table 11. For both RLHF-instructed and pre-trained models, the quality of output produced by the baselines tends to dramatically decrease at higher temperatures (only min-$p$ achieves good results at $\tau>1.0$), with the consequence of an increasing against-greedy diversity due to the choice of random and meaningless tokens. Instead, the quality of the output generated by DiffSampling remains more stable, with small but consistent increases in diversity.
F.3 Divergence Association Task
Figure 5 reports the DAT score and the percentage of output validity of the DPO-instructed and pre-trained models with different temperature values. Top-$p$ sampling, locally typical sampling, and $\eta$-sampling rapidly stop outputting valid lists of nouns when the temperature raises, even if the DAT score tends to be higher at $\tau=1.5$; min-$p$ returns a high percentage of valid outputs even at $\tau=2.0$, but increase the DAT score only for the pre-trained model and cannot produce anything valid at $\tau=10.0$. Instead, the performance of our methods remains very similar across different temperatures in terms of both the DAT score and the percentage of valid outputs, except for DiffSampling-lb, whose scores increase at the cost of some validity.
On the other hand, the greedy decoding strategy is less effective for the pre-trained model, which results in higher temperatures yielding better DAT scores across both the baselines and our methods. However, the number of valid outputs decreases faster, and top-$p$ sampling, locally typical sampling, and $\eta$-sampling produce very few correct lists at a temperature of $1.5$ (but with a higher DAT score). Again, min-$p$ better manages temperatures around $1.5$ and $2.0$, with higher scores and still at least half of the outputs as valid, but cannot produce any correct output at a temperature of $10.0$.
F.4 WritingPrompts
@@ F.5 Temperature Before or After Truncating
DiffS.-cut $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.62_{\pm.01}$ $0.41_{\pm.01}$ $0.37_{\pm.01}$ $0.64_{\pm.01}$ $0.36_{\pm.01}$ $0.23_{\pm.00}$ $0.49_{\pm.01}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$ $0.24_{\pm.01}$ $0.43_{\pm.01}$ $0.22_{\pm.01}$
DiffS.-lb $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.62_{\pm.01}$ $0.38_{\pm.01}$ $0.35_{\pm.01}$ $0.56_{\pm.01}$ $0.30_{\pm.01}$ $0.22_{\pm.00}$ $0.47_{\pm.01}$ $0.62_{\pm.01}$ $0.39_{\pm.01}$ $0.38_{\pm.01}$ $0.59_{\pm.01}$ $0.32_{\pm.01}$
DiffS.-minp $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.39_{\pm.01}$ $0.34_{\pm.01}$ $0.56_{\pm.01}$ $0.30_{\pm.01}$ $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.62_{\pm.01}$ $0.39_{\pm.01}$ $0.37_{\pm.01}$ $0.58_{\pm.01}$ $0.32_{\pm.01}$
@@
DiffS.-cut $0.23_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.37_{\pm.01}$ $0.28_{\pm.01}$ $0.50_{\pm.01}$ $0.26_{\pm.01}$ $0.23_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$ $0.25_{\pm.01}$ $0.46_{\pm.01}$ $0.23_{\pm.01}$
DiffS.-lb $0.06_{\pm.00}$ $0.16_{\pm.01}$ $0.27_{\pm.01}$ $0.64_{\pm.01}$ $0.73_{\pm.00}$ $0.89_{\pm.01}$ $0.79_{\pm.01}$ $0.20_{\pm.00}$ $0.43_{\pm.01}$ $0.57_{\pm.01}$ $0.40_{\pm.01}$ $0.52_{\pm.01}$ $0.71_{\pm.01}$ $0.45_{\pm.01}$
DiffS.-minp $0.20_{\pm.00}$ $0.44_{\pm.01}$ $0.58_{\pm.01}$ $0.38_{\pm.01}$ $0.51_{\pm.01}$ $0.71_{\pm.01}$ $0.45_{\pm.01}$ $0.21_{\pm.00}$ $0.46_{\pm.01}$ $0.60_{\pm.01}$ $0.35_{\pm.01}$ $0.45_{\pm.01}$ $0.67_{\pm.01}$ $0.41_{\pm.01}$
@@
DiffS.-cut $0.23_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.38_{\pm.01}$ $0.31_{\pm.01}$ $0.54_{\pm.01}$ $0.29_{\pm.01}$ $0.23_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$ $0.25_{\pm.01}$ $0.46_{\pm.01}$ $0.24_{\pm.01}$
DiffS.-lb $01_{\pm.00} 0.0$$03_{\pm.00}$ $0.13_{\pm.00}$ $5_{\pm.01} 0.66±.0$$6_{\pm.00}$ $0.92_{\pm.01}$ $0.92_{\pm.00}$ $0.19_{\pm.00}$ $0.43_{\pm.01}$ $0.56_{\pm.01}$ $0.43_{\pm.01}$ $0.54_{\pm.01}$ $0.73_{\pm.01}$ $0.48_{\pm.01}$
DiffS.-minp $0.19_{\pm.00}$ $0.42_{\pm.01}$ $0.54_{\pm.01}$ $0.42_{\pm.01}$ $0.57_{\pm.00}$ $0.76_{\pm.01}$ $0.50_{\pm.01}$ $0.21_{\pm.00}$ $0.46_{\pm.01}$ $0.59_{\pm.01}$ $0.37_{\pm.01}$ $0.47_{\pm.01}$ $0.69_{\pm.01}$ $0.42_{\pm.01}$
@@
DiffS.-cut $0.17_{\pm.00}$ $0.38_{\pm.01}$ $0.51_{\pm.01}$ $0.57_{\pm.01}$ $0.53_{\pm.01}$ $0.74_{\pm.01}$ $0.50_{\pm.01}$ $0.23_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$ $0.25_{\pm.01}$ $0.47_{\pm.01}$ $0.25_{\pm.01}$
DiffS.-lb $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.11_{\pm.00}$ $0.74_{\pm.01}$ $0.64_{\pm.00}$ $0.93_{\pm.00}$ $0.93_{\pm.00}$ $0.17_{\pm.00}$ $0.40_{\pm.01}$ $0.52_{\pm.01}$ $0.48_{\pm.01}$ $0.59_{\pm.01}$ $0.78_{\pm.01}$ $0.54_{\pm.01}$
DiffS.-minp $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.11_{\pm.00}$ $0.74_{\pm.01}$ $0.64_{\pm.00}$ $0.93_{\pm.00}$ $0.93_{\pm.00}$ $0.21_{\pm.00}$ $0.45_{\pm.01}$ $0.58_{\pm.01}$ $0.38_{\pm.01}$ $0.49_{\pm.01}$ $0.72_{\pm.01}$ $0.45_{\pm.01}$
@@
DiffS.-cut $0.21_{\pm.00}$ $0.48_{\pm.01}$ $0.70_{\pm.01}$ $0.50_{\pm.01}$ $0.29_{\pm.01}$ $0.47_{\pm.01}$ $0.27_{\pm.01}$ $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.00}$ $0.37_{\pm.01}$ $0.18_{\pm.00}$ $0.30_{\pm.01}$ $0.16_{\pm.01}$
DiffS.-lb $0.20_{\pm.00}$ $0.46_{\pm.01}$ $0.68_{\pm.01}$ $0.54_{\pm.01}$ $0.38_{\pm.01}$ $0.50_{\pm.01}$ $0.31_{\pm.01}$ $0.20_{\pm.00}$ $0.45_{\pm.01}$ $0.65_{\pm.01}$ $0.59_{\pm.01}$ $0.42_{\pm.01}$ $0.55_{\pm.01}$ $0.34_{\pm.01}$
DiffS.-minp $0.20_{\pm.00}$ $0.47_{\pm.01}$ $0.69_{\pm.01}$ $0.50_{\pm.01}$ $0.34_{\pm.01}$ $0.47_{\pm.01}$ $0.28_{\pm.01}$ $0.20_{\pm.00}$ $0.47_{\pm.01}$ $0.67_{\pm.01}$ $0.55_{\pm.01}$ $0.38_{\pm.01}$ $0.52_{\pm.01}$ $0.31_{\pm.01}$
@@
DiffS.-cut $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.01}$ $0.41_{\pm.01}$ $0.22_{\pm.00}$ $0.36_{\pm.01}$ $0.20_{\pm.01}$ $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.01}$ $0.38_{\pm.01}$ $0.19_{\pm.00}$ $0.33_{\pm.01}$ $0.18_{\pm.01}$
DiffS.-lb $03_{\pm.00} 0.0$$07_{\pm.00}$ $0.18_{\pm.00}$ $0.79_{\pm.01}$ $0.75_{\pm.00}$ $0.97_{\pm.00}$ $0.89_{\pm.00}$ $0.11_{\pm.00}$ $0.23_{\pm.01}$ $0.34_{\pm.01}$ $0.75_{\pm.01}$ $0.79_{\pm.00}$ $0.89_{\pm.00}$ $0.70_{\pm.01}$
DiffS.-minp $0.17_{\pm.00}$ $0.38_{\pm.01}$ $0.52_{\pm.01}$ $0.77_{\pm.01}$ $0.61_{\pm.00}$ $0.80_{\pm.01}$ $0.52_{\pm.01}$ $0.19_{\pm.00}$ $0.43_{\pm.01}$ $0.60_{\pm.01}$ $0.70_{\pm.01}$ $0.52_{\pm.00}$ $0.70_{\pm.01}$ $0.44_{\pm.01}$
@@
DiffS.-cut $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.72_{\pm.01}$ $0.44_{\pm.01}$ $0.24_{\pm.00}$ $0.39_{\pm.01}$ $0.22_{\pm.01}$ $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.01}$ $0.38_{\pm.01}$ $0.19_{\pm.00}$ $0.33_{\pm.01}$ $0.18_{\pm.01}$
DiffS.-lb $01_{\pm.00} 0.0$$03_{\pm.00}$ $0.14_{\pm.00}$ $0.80_{\pm.01}$ $0.66_{\pm.00}$ $0.98_{\pm.00}$ $94_{\pm.00} 0.0$$09_{\pm.00}$ $0.19_{\pm.01}$ $0.29_{\pm.01}$ $0.76_{\pm.01}$ $0.82_{\pm.00}$ $0.92_{\pm.00}$ $0.75_{\pm.01}$
DiffS.-minp $0.13_{\pm.00}$ $0.29_{\pm.01}$ $0.41_{\pm.01}$ $0.81_{\pm.01}$ $0.73_{\pm.00}$ $0.90_{\pm.00}$ $0.64_{\pm.01}$ $0.19_{\pm.00}$ $0.42_{\pm.01}$ $0.58_{\pm.01}$ $0.72_{\pm.01}$ $0.53_{\pm.00}$ $0.72_{\pm.01}$ $0.45_{\pm.01}$
@@
DiffS.-cut $0.07_{\pm.00}$ $0.17_{\pm.01}$ $0.32_{\pm.01}$ $0.53_{\pm.01}$ $0.58_{\pm.01}$ $0.73_{\pm.01}$ $0.71_{\pm.01}$ $0.21_{\pm.00}$ $0.49_{\pm.00}$ $0.73_{\pm.01}$ $0.39_{\pm.01}$ $0.19_{\pm.00}$ $0.34_{\pm.01}$ $0.19_{\pm.01}$
DiffS.-lb $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.75_{\pm.01}$ $0.63_{\pm.00}$ $0.97_{\pm.00}$ $95_{\pm.00} 0.0$$05_{\pm.00}$ $0.13_{\pm.00}$ $0.23_{\pm.01}$ $0.77_{\pm.01}$ $0.83_{\pm.00}$ $0.95_{\pm.00}$ $0.83_{\pm.01}$
DiffS.-minp $00_{\pm.00} 0.0$$02_{\pm.00}$ $0.12_{\pm.00}$ $0.75_{\pm.01}$ $0.63_{\pm.00}$ $0.97_{\pm.00}$ $0.95_{\pm.00}$ $0.18_{\pm.00}$ $0.40_{\pm.01}$ $0.55_{\pm.01}$ $0.76_{\pm.01}$ $0.57_{\pm.00}$ $0.79_{\pm.01}$ $0.50_{\pm.01}$
@@
F.6 Quality-Diversity Trade-Off Visualization
To simplify the comprehension of how temperature impacts the quality-diversity trade-off and how the temperature position changes the performance of different sampling methods, we present here the Pareto fronts for top-$p$, min-$p$, and our three DiffSampling methods (with temperature applied either before or after truncation) at five different temperatures $\tau$. We omit the DAT plots for conciseness, as they convey the same message as Figures 5 and 6.
Figure 7: Quality-diversity trade-offs for GSM8K and MATH datasets at five different temperatures for two baselines (top-$p$ and min-$p$) and our three DiffSampling methods, with temperature applied either before or after the truncation. Each point represents the mean score across the entire dataset and three different seeds.
Figure 7 reports the plots for the math problem-solving datasets by considering both against-greedy EAD diversity and against-greedy SBERT diversity with respect to accuracy. As already discussed, applying the temperature after preserves higher quality while reducing the increments in diversity. Vice versa, applying the temperature before has dramatic effects on accuracy, and only DiffSampling-cut achieves competitive results at very high $\tau$. In general, DiffSampling-minp with temperature applied after seems to provide the best balance between quality and diversity at different $\tau$.
Figure 8: Quality-diversity trade-offs for the XSum dataset with the instructed and pre-trained models at five different temperatures for two baselines (top-$p$ and min-$p$) and our three DiffSampling methods, with temperature applied either before or after the truncation. Each point represents the mean score across all the outputs generated from the same 1000 randomly sampled inputs.
The results for XSum are similar. As shown in Figure 8, the coherence of generated outputs degrades fast with $\tau>1$ if temperature is applied before truncation. While min-$p$ and DiffSampling-minp have interesting performances at $\tau\leq 2.0$, for very high temperatures only DiffSampling-cut achieves comparable coherence. Overall, all sampling methods apart from DiffSampling-cut have very similar quality-diversity trade-offs, but DiffSampling-minp with temperature applied after seems to provide the best results considering the full range of temperature values.
Figure 9: Quality-diversity trade-offs for the WritingPrompts dataset with the instructed and pre-trained models at five different temperatures for two baselines (top-$p$ and min-$p$) and our three DiffSampling methods, with temperature applied either before or after the truncation. Each point represents the mean score across all the outputs generated from the same 500 randomly sampled inputs.
Finally, Figure 9 shows the quality-diversity trade-offs in the case of the WritingPrompts dataset. Once more, DiffSampling-cut, especially with temperature applied after, better preserves quality at different $\tau$, while DiffSampling-minp with temperature applied after better balances quality and diversity for higher temperatures.
Appendix G Additional Experiments
G.1 Divergent Association Task over All Nouns
G.2 Ablation Study on the Lower Bound